What are the x- and y-intercepts of the line with equation 4x + 5y = 40?
Answer by PRMath(133)

4x + 5y = 40


Here are the two facts you need to know:


To find your X Intercept, solve for X by making y = 0.
To find your Y Intercept, solve for Y by making X = 0.

So let's do that. Let's do the y intercept first. I'll just "plug" 0 where "X" is in the equation:

4x + 5y = 40 (Original equation)
4(0) + 5y = 40 (Plugged 0 in for the X variable)
0 + 5y = 40
5y = 40
y = 8 (Divided both sides of the equation by 5 to isolate the "Y")
SO - now you know that when X = 0, then Y = 8. SO your y intercept is 8. You can use this ordered pair: (0, 8)


Now let's find the X intercept. I'll just "plug" 0 where "Y" is in the equation:
4x + 5y = 40 (Original equation)
4x + 5(0) = 40 (Plugged 0 in for the "Y" variable)
4x + 0 = 40
4x = 40
x = 10 (Divided both sides of the equation by 4 to isolate the "X")

SO - now you know that when Y = 0, then X = 10. So your X intercept is 10. You can use this ordered pair: (10, 0)

I hope this helps. :-)